Herschel Status Report - October 2011
19 Oct 2011 14:20
Report for period 22 September to 13 October 2011
Mission operations of the Herschel space observatory continued nominally during the reporting period, with the spacecraft and subsystems all performing as expected.
The spacecraft continues to be in good health and is operating nominally.
During the previous reporting period, on 18 September, a correction was implemented to correct for tilt of the CCD in the star tracker focal plane (see the previous status report linked from the right-hand menu). Following the successful trial, this correction was made permanent on 26 September.
While Herschel's pointing performance has always been well within specification, the correction has resulted in a further improvement of the Absolute Pointing Error (APE) of the observatory. The APE has been confirmed to now be about 1" (prior to the correction it was about 1.5"), corresponding to less than one-hundredth of the size of a CCD pixel.
As a result of this supreme pointing as compared to the specifications, higher order effects can now be seen. One follow-up that is being pursued consists of identifying and removing guide stars of lesser positional accuracy from the guide star catalogue that otherwise could lead to outliers (at the new, smaller scale) in the APE.
Operations for all three instruments, PACS, SPIRE and HIFI, have been nominal during the reporting period.
Ground segment operations continued to proceed well and the data recovery has remained at 100%. As of 7 October 2011, the approximate completion of the different programme parts was:
For more details of these different programme parts, see the "overview of Herschel observing" linked from the right-hand menu.
The second, and final, call for Open Time Proposals (OT2) for observing with Herschel was closed in mid-September. The technical evaluation of the 531 submitted proposals is in progress in preparation of the Herschel Observing Time Allocation Committee meeting in the first half of November.